• A wave has amplitude of 15 cm and displaces at a speed of 80 m/s at a frequency of 270 Hz

a) Calculate its wavelength


  • A longitudinal wave on a slinky toy has a frequency of 4Hz and a speed of 1.5 m/s. What is the wavelength of this wave?







1
Expert's answer
2021-09-09T17:21:25-0400

By definition, the wavelength is:


"\lambda = \dfrac{c}{\nu}"

where "c" is the speed of the wave, and "\nu" is its frequency.

a) In the first case obtain:


"\lambda = \dfrac{80m/s}{270Hz} \approx 0.296m"

b) In the second case obtain:


"\lambda = \dfrac{1.5m/s}{4Hz} = 0.375m"

Answer. a) 0.296 m, b) 0.375 m.